OpenStudy (anonymous):

What is the partial sum of ? a 2,272 b –2,208 c –2,272 d 2,208

OpenStudy (anonymous):

\[\sum_{i=10}^{25} -2-8i\]

OpenStudy (anonymous):

it is negative for sure, since all the terms are

OpenStudy (anonymous):

of course it is C answer is always C

OpenStudy (anonymous):

Can I have the real answer? or if this is the real answer can you show me how you got it? Thanks

OpenStudy (anonymous):

@Chlorophyll

OpenStudy (anonymous):

C is the real answer

OpenStudy (anonymous):

Okay, how do you get it?

OpenStudy (anonymous):

there is a snap way to do this, but i do not know it. i compute \[\sum_{i=1}^{25}-2-8i -\sum_{i=1}^9-2-8i\]

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

first one is easy enough, it is \[\sum_{i=1}^{25}-2-8\sum_{i=1}^{25}i\] \[=-50-8\times \frac{25\times 26}{2}=-50-4\times25 \times 26\] \[=-50-2600=-2650\] second part is similar

OpenStudy (anonymous):

That answer is not an option, but thank you anyway

OpenStudy (anonymous):

oh that was just the first part,

OpenStudy (anonymous):

Oh, alright

OpenStudy (anonymous):

you still have to compute \[\sum_{i=1}^9-2-8i\] and subtract it from the answer i wrote above. C is the final answer

OpenStudy (anonymous):

Thanks!